10. Eliquis The drug Eliquis (apixaban) is used to help prevent blood clots in certain patients. In clinical trials, among 5924 patients treated with Eliquis, 153 developed the adverse reaction of nausea (based on data from Bristol-Myers Squibb Co.). Use a 0.05 significance level to test the claim that 3% of Elíquis users develop nausea. Does nausea appear to be a problematic adverse reaction? irwOK a dultg or
Answer #1

H0: P = 0.03

H1: P \small \neq 0.03

\small \widehat p = 153/5924 = 0.0258

The test statistic z = (\small \widehat p - P)/sqrt(P(1 - P)/n)

                             = (0.0258 - 0.03)/sqrt(0.03 * 0.97/5924)

                             = -1.90

P-value = 2 * P(Z < -1.90)

             = 2 * 0.0287

             = 0.0574

Since the P-value is greater than the significance level(0.0574 > 0.05), so we shoould not reject the null hypothesis.

So there is sufficient evidence to support the claim that 3% of Eliquis users develop nausea.

Since the p-value is not less than 0.05, so it does not appear to be a problematic reaction.